# Donghyeon Lee1, Minbyul Jeong2∗, Jaewoo Kang1,3

### v2

@@ 1 Introduction
Figure 1: The overview of ChroKnowBench. We gather knowledge with time stamps and separate them in three key aspects: (1) multiple domains: general, biomedical, legal, commonsense, and mathematics; (2) time dependency: as time goes by, changeable knowledge or not; (3) temporal state: dynamic (has evolved over period) and static (no change occurred during period).
@@
To this end, our ChroKnowPrompt approach utilizes an in-depth chronological prompting strategy that traverses knowledge across adjacent time spans, effectively addressing issues of partial recall and temporal boundaries (Section 6). In knowledge recall, our evaluation reveals improvements in the biomedical domain (11.2%) and the general domain (2.6%), shifting knowledge category from Partial Known to Known for unchanged objects. Our non-parametric approach allows for direct updates to both proprietary and open-source LLMs without extensive retraining, highlighting practicality and broad applicability (Section 7). Our work emphasizes the importance of temporal context in eliciting LLMs knowledge while identifying challenges with dynamic datasets and unstructured, context-rich formats like in the legal domain. A comprehensive analysis advocates for integrating parametric techniques to complement prompting and achieve more robust temporal knowledge handling.
@@ 2.2 Eliciting Knowledge using diverse Templates
Since models prefer different formats when eliciting their knowledge, it is important to use varied approaches to accurately assess their understanding (Zhou et al., ). While we initially evaluate the model’s knowledge using a standard triplet format, relying on a single template may not sufficiently capture the full extent of the model’s knowledge. Thus, following Hendrycks et al. (2021); Huang et al. (2024), we also employ a well-known format, multiple-choice question answering (MCQA) with 4 options, and True/False to elicit the model’s knowledge more effectively. As a result, we propose three templates for measuring how much knowledge the model holds: triplets (hereafter referred to as Generation), MCQA, and TF. Each template is designed with appropriate few-shot exemplars and corresponding matching rules. For example, in Generation, due to the complexity of evaluating responses, we apply fuzzy matching techniques to compare the generated responses against predefined labels. See Appendix A.2 for further details of few-shot exemplars, fuzzy matching rules, and examples of three templates.
@@ 4 Experimental Setup
We conduct the experiments on a variety of proprietary and open-source LLMs. Each model utilizes either an instruction-tuned or chat version to enhance instruction following during sampling. We enumerate the eight open-source and two proprietary LLMs: Llama-3.1-8B-Instruct (Meta, 2024), Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7b-chat-hf (Touvron et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3.5-mini-instruct (Abdin et al., 2024), SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023), gemma-2-9b-it (Team et al., 2024b), gemma-7b-it (Team et al., 2024a) for major open-source models, and GPT-4o mini (OpenAI, 2024a), Gemini-1.5-flash (DeepMind, 2024) for proprietary models. We focus on anlayzing trends in the chronological knowledge captured by those models, differ in corpus coverage. Details of our inference setups are in Appendix A.4.
@@ 7.1 Results of ChroKnowPrompt
Details of task configuration is in Appendix A.8. Figure 7 presents the effect of chronological prompting on unchanged objects across different models. Results are shown as the rise of percentage in Known category with decreasing Partial Known, indicating the increase by Chrono-correct. Significant overall improvements are observed in the biomedical domain (average increase of 11.2%), while general and legal domains show smaller gains (2.6% and 3.2%, respectively). This indicates that chronological prompting effectively enhances knowledge recall without requiring external retrieval systems. In the legal domain, models struggle to recall knowledge due to the complexity of unstructured, context-rich data, especially for dynamic dataset. Among the models, proprietary GPT4o-mini and Gemini-1.5-flash perform well, while open-source models like Llama-2-7b-chat-hf also show notable improvements despite outdated.
@@ 7.2 Analysis
Table 5 compares results for changed and unchanged objects in dynamic datasets across domains. As shown in Section 7.1, ChroKnowPrompt effectively aids models in recalling unchanged objects in both dynamic and static datasets. However, its performance on changed objects in dynamic datasets remains limited, achieving only 10–30% performance of unchanged object cases (an average of 0.43 in general domains). This highlights, despite the helpfulness and applicability of ChroKnowPrompt in addressing the chronological gaps of knowledge, recalling all historical changes of objects using prompting alone remains an exceptionally challenging problem, particularly in complex contexts such as the legal domain. These findings emphasize the need for further research into effective parametric editing techniques to assist temporal knowledge handling.
@@
To elucidate the mechanisms of chronological prompting, we analyze the impact of incorporating the next span in chronological contexts. As shown in Table 6, the total span (both previous and next) yields higher scores than using only the previous span with the degree of improvement varying by domain. In the biomedical domain, the total span nearly doubles the score of the previous span alone (11.9 vs. 6.1), while the general domain shows a modest increase (2.8 vs. 1.8). Model-specific temporal sensitivity also varies: Llama-2-7b-chat-hf effectively utilizes next spans, whereas Gemini-1.5-flash and SOLAR-10.7B-Instruct-v1.0 benefit more from previous spans. These suggest that differences in temporal context utilization and the coverage of pretraining corpus may influence models’ sensitivity and knowledge recall across time frames.
@@
Additionally, we analyze various chat models (before instruction-tuning), including Llama-2-7b-chat-hf with findings that chronological prompting can enhance both current and legacy models by leveraging temporal context. We evaluate three open-source chat models: mpt-7b-chat (Research, 2023), Pythia-Chat-Base-7B (Biderman et al., 2023), and nemotron-3-8b-chat-4k-sft-hf (Zhang et al., 2023a). As shown in Table 6 and 7, ChroKnowPrompt is not particularly effective for chat models. Only mpt-7b-chat achieves a comparable peformance to Llama-2-7b-chat-hf (an average increase of 11.4), while Pythia-Chat-Base-7B shows almost no improvement.
@@ 8 Related work
Since the emergence of LMs, deriving knowledge from language model is extensively studied, such as probing tasks (Hewitt & Manning, 2019), LAMA (Petroni et al., 2019) and BioLAMA (Sung et al., 2021). Then, many subsequent studies follows to exploit, (1) how LLMs define knowledge (Yu et al., ; Zhang et al., 2023b; Gottesman & Geva, 2024), (2) how these models represent it (Chen et al., 2024a; b; Wang et al., 2024d), and (3) how manipulate misleading part (Wang et al., 2023; Gutiérrez et al., 2024; Wu et al., 2024a). Based on them, recent investigations of knowledge highlight the dynamic nature of evolving facts and suggests that contradictions within the training data may lead to knowledge conflicts (Marjanović et al., 2024; Chang et al., 2024; Wang et al., 2024a; Xu et al., 2024; Jin et al., 2024). And knowledge overshadowing (Zhang et al., 2024c) reveals phenomena where certain conditions overshadow other facts, leading to misleading information (i.e., hallucinations).
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea [NRF-2023R1A2C3004176, RS-2023-00262002], the Ministry of Health & Welfare, Republic of Korea [HR20C002103], and the ICT Creative Consilience program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the MSIT [IITP-2024-2020-0-01819].
@@ A.1.1 Parametric Knowledge Update
Considering update of LLMs are in two types, parametric and non-parametric (Wang et al., 2024c), a classical way of parametric update is using fine-tuning (Ghosal et al., 2024; Mecklenburg et al., 2024; Ge et al., 2024). While the method extended to LoRA (Hu et al., ), QLoRA (Dettmers et al., ), Melo (Yu et al., 2024) to continual learning (GRACE (Hartvigsen et al., 2024) and WISE (Wang et al., 2024b)), by using the parameter accessibility of white-box LLMs like Llama series (Touvron et al., 2023a), MEND (Mitchell et al., ), ROME (Meng et al., 2022), MEMIT (Meng et al., ) emerges. Those local editable methods are effective, and still try to improve specificity and generalizability.
@@ A.1.2 Non-parametric Knowledge Update
In contrast, for black-box LLMs, updates rely on non-parametric knowledge methods (Onoe et al., 2023), such as SERAC (Mitchell et al., 2022), MeLLo (Zhong et al., 2023), and IKE (Zheng et al., 2023). They align with two key trends: (1) Mitigating catastrophic forgetting, where the model loses previous knowledge, by not directly updating parameters. (2) Exploiting abilities of prominent black-box LLMs like GPT-o1 (OpenAI, 2024b) and Gemini (Team et al., 2023), as we cannot access to parameters. Another concern in knowledge update is they often focus only structured format, pointed out that current methods struggle to update unstrctured data effectively (Wu et al., 2024b). In this paper, we focus on non-parametric knowledge updates accomodating a broad range of input formats (structured and unstructured) to represent knowledge across diverse domains, depending on the use of various white-box and black-box LLMs.
@@ A.6 Total Time Frame Result of ChroKnowledge
Figure 10 represents the total results for all the range of time frames in general, and legal domain. Figure 11 to Figure 15 represents the total results of template-wise performance in ChroKnowledge. Each domain’s result is separated into two category of temporal state: Dynamic and Static. Every spider plots consist with three template: Generation, Multi-choice QA, and True/False. Each statistics refer the percentage of Correct answer, same as Figure 2.
@@ A.8 Task Configurations of ChroKnowPrompt
Verify & Refine Case
@@
Object Increase 0.4 2.9 0.9 11.8 0.0 3.2
Temporal Increase 1.7 2.6 5.9 12.7 0.2 5.8
Domain Increase 2.0 8.2 2.1
@@ A.9 Details in span-wise results of ChroKnowPrompt
Temporal Increase 3.2 2.5 2.3 1.4 11.5 12.7 6.6 6.8
Domain Increase 2.9 1.8 12.1 6.8
@@
Temporal Increase 0.4 5.0 0.3 3.8
Domain Increase 2.7 2.1
Figure 10: Total result of performance heatmap in general and legal domain

### v3

# Donghyeon Lee1,3, Minbyul Jeong2∗, Jaewoo Kang1,3

@@ 1 Introduction
Figure 1: The overview of ChroKnowBench. We gather knowledge with time stamps and separate them in three key aspects: (1) multiple domains: general, biomedical, legal, commonsense, and mathematics; (2) time dependency: as time goes by, changeable knowledge or not; (3) temporal state: dynamic (has evolved over period) and static (no change occurred during period). Here, trends of Correct2.1) for each years represented by line plots show difference among domains and temporal states. And each highlighted portions are chronologically Known following §6.1.
@@
To this end, our ChroKnowPrompt approach utilizes an in-depth chronological prompting strategy that traverses knowledge across adjacent time spans, effectively addressing issues of partial recall and temporal boundaries (Section 6). In knowledge recall, our evaluation reveals improvements in the biomedical domain (11.5%) and the general domain (2.8%), shifting knowledge category from Partial Known to Known for unchanged objects. Our non-parametric approach allows for direct updates to both proprietary and open-source LLMs without extensive retraining, highlighting practicality and broad applicability (Section 7). Our work emphasizes the importance of temporal context in eliciting LLMs knowledge while identifying challenges with dynamic datasets and unstructured, context-rich formats like in the legal domain. A comprehensive analysis advocates for integrating parametric techniques to complement prompting and achieve more robust temporal knowledge handling.
@@ 2.2 Eliciting Knowledge using diverse Templates
Since models prefer different formats when eliciting their knowledge, it is important to use varied approaches to accurately assess their understanding (Zhou et al., 2024). While we initially evaluate the model’s knowledge using a standard triplet format, relying on a single template may not sufficiently capture the full extent of the model’s knowledge. Thus, following Hendrycks et al. (2021); Huang et al. (2024), we also employ a well-known format, multiple-choice question answering (MCQA) with 4 options, and True/False to elicit the model’s knowledge more effectively. As a result, we propose three templates for measuring how much knowledge the model holds: triplets (hereafter referred to as Generation), MCQA, and TF. Each template is designed with appropriate few-shot exemplars and corresponding matching rules. For example, in Generation, due to the complexity of evaluating responses, we apply fuzzy matching techniques to compare the generated responses against predefined labels. See Appendix A.2 for further details of few-shot exemplars, fuzzy matching rules, and examples of three templates.
@@ 4 Experimental Setup
We enumerate the nine open-source and two proprietary LLMs: Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct (Meta, 2024), Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7b-chat-hf (Touvron et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3.5-mini-instruct (Abdin et al., 2024), SOLAR-10.7B-Instruct-v1.0 (Kim et al., 2023), gemma-2-9b-it (Team et al., 2024b), gemma-7b-it (Team et al., 2024a) for major open-source models, and GPT-4o mini (OpenAI, 2024a), Gemini-1.5-flash (DeepMind, 2024) for proprietary models. Each model utilizes either an instruction-tuned or chat version to enhance instruction following during sampling. We focus on anlayzing trends in the chronological knowledge captured by those models, differ in corpus coverage. Details of our inference setups are in Appendix A.4.
@@ 7.1 Results of ChroKnowPrompt
Details of task configuration is in Appendix A.8. Figure 7 presents the effect of chronological prompting on unchanged objects across different models. Results show the rise of percentage in Known category with decreasing Partial Known, indicating the increase by Chrono-correct. Significant improvements are observed in the biomedical domain (average increase of 11.5%), while general and legal domains show smaller gains (2.8% and 3.1%, respectively). Both proprietary models (GPT4o-mini, Gemini-1.5-flash) and the massive open-source model (Llama-3.1-70B-Instruct) perform well, while smaller open-source models like Llama-2-7B-chat-hf also show notable improvements despite being outdated. This indicates that chronological prompting effectively enhances knowledge recall without requiring external retrieval systems. But in the legal domain, models struggle to recall knowledge due to the complexity of unstructured, context-rich data, especially for dynamic dataset.
@@ 7.2 Analysis
Table 5 compares results for changed and unchanged objects in dynamic datasets across domains. As shown in Section 7.1, ChroKnowPrompt effectively aids models in recalling unchanged objects in both dynamic and static datasets. However, its performance on changed objects in dynamic datasets remains limited, achieving only 10–30% performance of unchanged object cases (an average of 0.4 in general domains). This highlights, despite the helpfulness and applicability of ChroKnowPrompt in addressing the chronological gaps of knowledge, recalling all historical changes of objects using prompting alone remains an exceptionally challenging problem, particularly in complex contexts such as the legal domain. These findings emphasize the need for further research into effective parametric editing techniques to assist temporal knowledge handling.
@@
To elucidate the mechanisms of chronological prompting, we analyze the impact of incorporating the next span in chronological contexts. As shown in Table 6, the total span (both previous and next) yields higher scores than using only the previous span with the degree of improvement varying by domain. In the biomedical domain, the total span nearly doubles the score of the previous span alone (12.0 vs. 6.7), while the general domain shows a modest increase (2.8 vs. 1.8). Model-specific temporal sensitivity also varies: Llama-2-7b-chat-hf effectively utilizes next spans, whereas Gemini-1.5-flash and SOLAR-10.7B-Instruct-v1.0 benefit more from previous spans. These suggest that differences in temporal context utilization and the coverage of pretraining corpus may influence models’ sensitivity and knowledge recall across time frames.
@@
Additionally, we analyze various chat models (before instruction-tuning), including Llama-2-7b-chat-hf, as chronological prompting may enhance both current and legacy models by leveraging temporal context. We evaluate three open-source chat models: mpt-7b-chat (Research, 2023), Pythia-Chat-Base-7B (Biderman et al., 2023), and nemotron-3-8b-chat-4k-sft-hf (Zhang et al., 2023a). As shown in Table 6 and 7, ChroKnowPrompt is not particularly effective for chat models. Only mpt-7b-chat achieves a comparable peformance to Llama-2-7b-chat-hf (an average increase of 11.4), while Pythia-Chat-Base-7B shows almost no improvement.
@@ 8 Related work
Since the emergence of LMs, deriving knowledge from language model is extensively studied, such as probing tasks (Hewitt & Manning, 2019), LAMA (Petroni et al., 2019) and BioLAMA (Sung et al., 2021). Then, many subsequent studies follows to exploit, (1) how LLMs define knowledge (Yu et al., 2023; Zhang et al., 2023b; Gottesman & Geva, 2024), (2) how these models represent it (Chen et al., 2024a; b; Wang et al., 2024d), and (3) how manipulate misleading part (Wang et al., 2023; Gutiérrez et al., 2024; Wu et al., 2024a). Based on them, recent investigations of knowledge highlight the dynamic nature of evolving facts and suggests that contradictions within the training data may lead to knowledge conflicts (Marjanović et al., 2024; Chang et al., 2024; Wang et al., 2024a; Xu et al., 2024; Jin et al., 2024). And knowledge overshadowing (Zhang et al., 2024c) reveals phenomena where certain conditions overshadow other facts, leading to misleading information (i.e., hallucinations).
@@ Acknowledgments
This work was supported in part by the National Research Foundation of Korea [NRF-2023R1A2C3004176, RS-2023-00262002], the Ministry of Health & Welfare, Republic of Korea [HR20C002103], and the ICT Creative Consilience program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the MSIT [IITP-2025-2020-0-01819].
@@ A.1.1 Parametric Knowledge Update
Considering update of LLMs are in two types, parametric and non-parametric (Wang et al., 2024c), a classical way of parametric update is using fine-tuning (Ghosal et al., 2024; Mecklenburg et al., 2024; Ge et al., 2024). While the method extends to LoRA (Hu et al., 2022), QLoRA (Dettmers et al., 2023), and Melo (Yu et al., 2024), as well as continual learning approaches such as GRACE (Hartvigsen et al., 2024) and WISE (Wang et al., 2024b), the parameter accessibility of open-source LLMs like the Llama series (Touvron et al., 2023a) enables techniques such as MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023) to emerge. Those local editable methods are effective, and still try to improve specificity and generalizability.
@@ A.1.2 Non-parametric Knowledge Update
In contrast, for black-box LLMs, updates rely on non-parametric knowledge methods (Onoe et al., 2023), such as SERAC (Mitchell et al., 2022b), MeLLo (Zhong et al., 2023), and IKE (Zheng et al., 2023). They align with two key trends: (1) Mitigating catastrophic forgetting, where the model loses previous knowledge, by not directly updating parameters. (2) Exploiting abilities of prominent black-box LLMs like GPT-o1 (OpenAI, 2024b) and Gemini (Team et al., 2023), as we cannot access to parameters. Another concern in knowledge update is they often focus only structured format, pointed out that current methods struggle to update unstrctured data effectively (Wu et al., 2024b). In this paper, we focus on non-parametric knowledge updates accomodating a broad range of input formats (structured and unstructured) to represent knowledge across diverse domains, depending on the use of various white-box and black-box LLMs.
@@ A.6 Total Time Frame Result of ChroKnowledge
Figure 1012 represent the total results with total time frames in general, biomedical and legal domain, including chat template models like mpt-7B. Figure 13– 17 represents the total results of template-wise performance in ChroKnowledge. Each domain’s result is separated into two temporal state: Dynamic and Static. Every spider plots consist with three template: Generation, Multi-choice QA, and True/False. Each statistics refer the percentage of Correct answer, same as Figure 2.
@@ A.8 Task Configurations of ChroKnowPrompt
Verification & Refinement Case
@@
Object Increase 0.4 2.8 1.0 11.6 0.0 3.1
Temporal Increase 1.7 2.6 6.0 12.3 0.3 5.7
Domain Increase 2.0 8.1 2.1
@@ A.9 Details in span-wise results of ChroKnowPrompt
Temporal Increase 3.1 2.5 2.3 1.4 11.6 12.4 6.7 6.6
Domain Increase 2.8 1.8 12.0 6.7
@@
Temporal Increase 0.5 4.9 0.3 3.7
Domain Increase 2.7 2.0
Figure 10: Total result of performance heatmap in general domain for all models
Figure 11: Total result of performance heatmap in biomedical domain for all models
Figure 12: Total result of performance heatmap in legal domain for all models